Solve this problem.8(3n+5)=-32
Answer by TimothyLamb(4379)

8(3n+5)=-32
24n+40=-32
24n=-32-40
n=(-32-40)/24
n=-3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php